Two dairy farmers produce milk for a local town with local milk demand given by P = 300 - 3Q (P denotes price measured in Rands, Q denotes the quantity measured in liters). Both farmers have the same cost function given by TC = 150 + 2Q (where denotes output).


(a) Suppose that both farmers decide to form a cartel, determine profits for each farmer under the cartel     


(b) What output should farmer 1 produce if he/she expects their rival to produce 20 units?  



1
Expert's answer
2021-06-28T17:20:03-0400

"TR = P\u00d7Q"

"TR = (300-3Q)Q"

"TR = 300Q -3Q^2"

"MR = \frac{dTR}{dQ}"


"MR = 300-6Q"

"MC = \frac{dTC}{dQ}"

"MC = 0 + 2"

"= 2"

"MC = MR"

"2 = 300-6Q"

"6Q = 300 - 2"

"Q = 298\/6"

"Q = 49.67"

"P = 300 -3(49.67)"

"P = 150.99"

"= \frac{150.99}{2}"

"Profit for each = 75.495"